# Digital Twin Vehicular Edge Computing Network: Task Offloading and Resource Allocation

### v1

@@ I Introduction
With the development of Internet of Things, more and more in car applications are being installed in modern cars[1]-[10]. However, due to the insufficient computing capability of the vehicle itself, it may not be possible to handle the computing tasks generated by these vehicular applications in a timely manner[11]-[17]. Vehicular edge computing (VEC), as a promising technology, can offload computing tasks to the VEC server or roadside unit, and assist in processing computing tasks by using the computing resources of the VEC server itself[18]-[28]. However, in such a way, VEC also faces new issues such as vehicle mobility and environmental dynamics.
Digital twin, as an emerging technology, can bridge the gap between entities in physical space and digital space of information mapping[29]-[31]. Based on historical data and real-time status, establishing digital replicas for physical entities can enhance the real-time interaction, realize close monitoring, by reliable communication between digital space and physical systems[32].
In terms of vehicle task processing, applying digital twin technology to VEC task processing scenarios can better monitor vehicle operation status to control and or change it.In [33], Dai et al. proposed an adaptive digital twin VEC network that utilizes digital twin caching content and designs an offloading scheme based on the DRL framework to minimize offloading latency. In [1], Cao et al. applied digital twins to the task offloading scenario of VEC and solved the vehicle mobility problem by constructing a multiple objective joint optimization problem. In [34], Tan et al. applied modular processing of digital twins to the Internet of Vehicles, improving task scheduling efficiency and reducing energy consumption. However, they only considered the situation where the vehicle only generates one task in a single time slot. In practice, vehicles may have multiple demands at the same time, so it is possible to have multiple tasks in one time slot.
@@ II System model
As shown in Figure 1, we consider a DT scenario of VEC network in which a vehicle generates multiple tasks in a single time slot. It consists of two layers, namely the physical layer and the cloud digital twin layer. The physical layer includes $N$ vehicles ($N\stackrel{{\scriptstyle\Delta}}{{=}}\{1,2,3,...,N\}$) and a single base station equipped with edge computing servers. The cloud digital twin layer is the digital mapping of each vehicle on the cloud. Specifically, each vehicle will generate $K$ tasks at a specific time slot. Due to the limitations of the vehicle’s own computing resources, the digital twin in the cloud will make offloading decisions for the generated $K$ tasks to determine whether they are to be processed locally or offloaded. If they are to be offloaded to the VEC server for processing, the digital twin will also allocate computing resources for the processing of each task. It should be noted that we assume that the next task will only be generated after all K tasks for a single vehicle have been completed.
@@ III-A Problem Formulation
Due to the mutual influence of resource allocation among tasks in problem P1, which includes integer and real components that are coupled with each other, it is an NP hard problem that cannot be solved using conventional methods in polynomial time. Therefore, we use multi-agent methods to transform the problem. For multi-agent MDP, we use a tuple $(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{R})$ to represent it, where $\mathcal{N}$ represents the set of agents, $\mathcal{S}$ represents the state space, $\mathcal{A}$ represents the action space, and $\mathcal{R}$ represents the reward function.
@@ III-B Solution
$\theta_{Q_{n}}^{{}^{\prime}}=\eta\theta_{Q_{n}}+(1-\eta)\theta_{Q_{n}}^{{}^{\prime}},$ (19)
$\theta_{\pi_{n}}^{{}^{\prime}}=\eta\theta_{\pi_{n}}+(1-\eta)\theta_{\pi_{n}}^{{}^{\prime}},$ (20)
where $\eta$ is parameter update rate, which $\eta\in[0,1]$. $\theta_{\pi_{n}}^{{}^{\prime}}$ and $\theta_{\pi_{n}}$ are the parameters for estimation actor network and target actor network, respectively.

### v2

@@ I Introduction
With the development of Internet of Things, more and more in car applications are being installed in modern cars[1]. However, due to the insufficient computing capability of the vehicle itself, it may not be possible to handle the computing tasks generated by these vehicular applications in a timely manner. Vehicular edge computing (VEC), as a promising technology, can offload computing tasks to the VEC server or roadside unit, and assist in processing computing tasks by using the computing resources of the VEC server itself[2][3]. However, in such a way, VEC also faces new issues such as vehicle mobility and environmental dynamics.
Digital twin, as an emerging technology, can bridge the gap between entities in physical space and digital space of information mapping[4],[5],[6]. Based on historical data and real-time status, establishing digital replicas for physical entities can enhance the real-time interaction, realize close monitoring, by reliable communication between digital space and physical systems[7].
In terms of vehicle task processing, applying digital twin technology to VEC task processing scenarios can better monitor vehicle operation status to control and or change it.In [8], Dai et al. proposed an adaptive digital twin VEC network that utilizes digital twin caching content and designs an offloading scheme based on the DRL framework to minimize offloading latency. In [1], Cao et al. applied digital twins to the task offloading scenario of VEC and solved the vehicle mobility problem by constructing a multiple objective joint optimization problem. In [9], Hevesli et al. utilized DT to predict the real-time status of IoT devices in edge air-ground network for more efficient resource allocation. In [10], Tan et al. applied modular processing of digital twins to the Internet of Vehicles, improving task scheduling efficiency and reducing energy consumption. However, they only considered the situation where the vehicle only generates one task in a single time slot. In practice, vehicles may have multiple demands at the same time, so it is possible to have multiple tasks in one time slot.
@@ II System model
As shown in Figure 1, we consider a DT scenario of VEC network in which a vehicle generates multiple tasks in a single time slot. It consists of two layers, namely the physical layer and the cloud digital twin layer. The physical layer includes $\mathbf{N}$ vehicles ($\mathbf{N}\stackrel{{\scriptstyle\Delta}}{{=}}\{1,2,3,...,N\}$) and a single base station equipped with edge computing servers. The cloud digital twin layer is the digital mapping of each vehicle on the cloud. Specifically, each vehicle will generate $K$ tasks at a specific time slot. Due to the limitations of the vehicle’s own computing resources, the digital twin in the cloud will make offloading decisions for the generated $K$ tasks to determine whether they are to be processed locally or offloaded. If they are to be offloaded to the VEC server for processing, the digital twin will also allocate computing resources for the processing of each task. It should be noted that we assume that the next task will only be generated after all K tasks for a single vehicle have been completed.
@@ III-A Problem Formulation
Due to the mutual influence of resource allocation among tasks in problem P1, which cannot be solved using conventional methods in polynomial time. Therefore, we use multi-agent methods to transform the problem. For multi-agent MDP, we use a tuple $(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{R})$ to represent it, where $\mathcal{N}$ represents the set of agents, $\mathcal{S}$ represents the state space, $\mathcal{A}$ represents the action space, and $\mathcal{R}$ represents the reward function.
@@ III-B Solution
$\theta_{Q_{n}}^{{}^{\prime}}=\epsilon\theta_{Q_{n}}+(1-\epsilon)\theta_{Q_{n}}^{{}^{\prime}},$ (19)
$\theta_{\pi_{n}}^{{}^{\prime}}=\epsilon\theta_{\pi_{n}}+(1-\epsilon)\theta_{\pi_{n}}^{{}^{\prime}},$ (20)
where $\epsilon$ is parameter update rate, which $\epsilon\in[0,1]$. $\theta_{\pi_{n}}^{{}^{\prime}}$ and $\theta_{\pi_{n}}$ are the parameters for estimation actor network and target actor network, respectively.